RC meeting and Run Plan, Monday March 13th
Return to the RC Meeting Calendar
- RC: Holly
- RC Daily report: https://logbooks.jlab.org/entry/4148882
We ran on 35uA throughout the day shift yesterday. We had a strange incident during the NMR sweep at 12pm where the sweep stopped in the middle with a message about a memory problem (https://logbooks.jlab.org/entry/4148271). The second attempt to sweep after this yielded a rather lower polarization (~37%). Bill later noted that it looks like there is a blip where the Helmholtz coils dropped to 0 for a moment and could be responsible for the loss of polarization. After this, Bill raised the over temperature (https://logbooks.jlab.org/entry/4148278) and restarted 2 laser fibers (https://logbooks/entry/4148279). Mid-way through the owl shift, we experienced an increase in oven temperature and a simultaneous decrease in the cell temperature (https://logbooks.jlab.org/entry/4148708). The next 4 subsequent NMR sweeps showed loss of polarization and no signal by the end (see logbook entries https://logbooks.jlab.org/entry/4148668, https://logbooks.jlab.org/entry/4148673, https://logbooks.jlab.org/entry/4148683, https://logbooks.jlab.org/entry/4148698). We took harp scans and checked for the carbon hole, but it's gone, see https://logbooks.jlab.org/entry/4148815. We officially lost the cell. See https://logbooks.jlab.org/entry/4148858
Owl summary: https://logbooks.jlab.org/entry/4148610 Swing summary: https://logbooks.jlab.org/entry/4148570
- While we request and are able to run at 45uA, accelerator has us running at 35uA for stability with Hall C online.
- Latest carbon hole image: https://logbooks.jlab.org/entry/4147220
- Running 45uA on pol-3He
- Production on 3He!
- GENII-3Stream config (currently no SBSGEMs until we can put them back in)
- PS1=2, PS3=0, PS5=6
- Useful guide from Bill on rates to compare when bringing beam back: https://logbooks.jlab.org/entry/4144886
Data Collection and Polarization Status
- Currently 0%: https://logbooks.jlab.org/entry/4148698, last best was 42.8%: https://logbooks.jlab.org/entry/4148551
- Data status at 4th-pass: 51.3 C (uncorrected)